# The Impact of Baryonic Effects on the Dynamical Masses inferred using Satellite Kinematics

### v1

@@ 4.2 Comparison with EAGLE
Fig. 3 compares the fiducial model to halos in EAGLE for two different masses: $M_{\rm DMO}=10^{12.5}\>{\rm M_{\odot}}$ (left panel) and $M_{\rm DMO}=10^{13.5}\>{\rm M_{\odot}}$ (right panel). Solid lines show the components of the analytical model, with different colors corresponding to different mass components, as indicated. To quantify the halo-to-halo variation in EAGLE halos, we compute the central 95% spread in density in bins of $r/r_{200}$ for two stacks of EAGLE halos whose median DMO masses match the model: $10^{12.25}<M_{\rm DMO}/\>{\rm M_{\odot}}<10^{12.85}$ (480 halos, left panel) and $10^{13.3}<M_{\rm DMO}/\>{\rm M_{\odot}}<10^{13.9}$ (31 halos, right panel). This spread is shown as shaded regions in matching colors for each component.
@@
Finally, the total density profiles show good agreement between the EAGLE simulations and our fiducial analytical model. This is noteworthy because although the simulation data was used to calibrate $f_{\rm eject}(M_{\rm DMO})$, the model was not tuned to reproduce any of the density profiles in EAGLE. Since the kinematics of the satellite galaxies are ultimately controlled by the total enclosed mass profile, we can thus be confident that our fiducial analytical model is meaningful.
@@ 4.3 Variations from the fiducial model
Figure 4: Left: Solving for the response of the dark matter due to the presence of the baryons for the fiducial model with $\nu=1$ for different halo masses. The final radius $r_{\rm f}$ is plotted against initial radius $r_{\rm i}$. The dashed line shows $r_{\rm f}=r_{\rm i}$, indicating when there would be no halo response. For small $r_{\rm i}$ we see that $r_{\rm f}<r_{\rm i}$, indicating that dark matter that was initially at larger radii migrated inwards (contraction). On the other hand, at at large $r_{\rm i}$, we see that $r_{\rm f}>r_{\rm i}$, indicating expansion. Both effects become weaker with increasing halo mass. Middle: The line-of-sight velocity dispersion profile for a halo of mass $M_{\rm DMO}\sim 10^{12.5}\>{\rm M_{\odot}}$, for $\nu=0$ (blue), $\nu=0.5$ (red) and $\nu=1$ (green). Under $\nu>0$, we again see that the dark matter halo responds by contracting in the center and expanding in the outskirts, though only the latter affects satellite kinematics data (gray shaded region). Right: Impact of the halo response on the aperture velocity dispersion ratio, $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$, as a function of halo mass. The blue curve indicates no halo response $\nu=0$, equivalent to $r_{\rm i}=r_{\rm f}$. When the dark matter responds ($\nu>0$), the observed effect is a further decrease of $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$.
@@ 4.3.1 Varying halo response strength
Figure 5: Top panel: The cumulative mass fractions of each baryonic component as a function of halo mass for the three scenarios considered (see text): the ’no-ejection model’ (left), the fiducial model (middle), and the ’extreme-ejection model’ (right). In each scenario, we keep the mass fractions of the central galaxy (red) and diffuse stellar component (pink) fixed, and we only vary the parameter $f_{\rm eject}$ (brown), which leads to different total gas fractions (blue). The total bound baryonic fraction, shown as the black dashed line, is simply obtained by summing all bound baryonic fractions, which is by definition equal to 1 minus the ejected fraction. Bottom panel: the impact of baryonic effects on the ratio of integrated velocity dispersion $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$ as a function of halo mass. Without baryonic corrections, the ratio remains at 1 (dotted black line). The blue curves, assuming no dark matter halo response, follow $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}\approx$ $\sqrt{M_{\rm tot}/M_{\rm DMO}}=\sqrt{1-f_{\rm b}f_{\rm eject}}$ (dashed black line). The green curves include halo response when preserving adiabatic invariants ($\nu=1$), further reducing velocity dispersion ratio. The gray shaded region represents the range defined by the extreme models, spanning zero to maximal baryonic effect.
@@ 4.3.2 Varying ejection strength
In this section, we examine the impact of changes in $f_{\rm eject}(M_{\rm DMO})$. Since mass conservation requires $f_{\rm gas}+f_{\rm*,cen}+f_{\rm*,diffuse}+f_{\rm eject}=1$, such changes also imply adjustment to one or more other mass fractions. To fully understand the consequences of these trade-offs, we test multiple configurations and define two limiting models that represent extreme, yet self-consistent, baryonic configurations.
@@
In the first extreme scenario, we keep $f_{\rm*,cen}$ and $f_{\rm*,diffuse}$ at their fiducial values, but set $f_{\rm eject}=0$. This means that we are not ejecting any baryons; for halos at all mass scales, the gas now complements the stars such that the total bound baryon fraction adds up to the universal baryon fraction. Since the gas contributes to the mass in the outskirts of the halo, where it has a density profile that is similar to that of the dark matter, this model leads to minimal baryonic effects on satellite kinematics. We refer to this configuration as the ‘no-ejection’ model.
@@
In the second extreme model, we again keep the fiducial fractions for the central galaxy and diffuse stellar component, but we change the gas ejection parameters that describe the mass dependence of $f_{\rm eject}$ to $\alpha=1.2$, $\log M_{\mathrm{char}}=13.7$ and $f_{\mathrm{eject,max}}=0.8$. These values give $f_{\rm eject}\sim 0.8$ at $10^{12}M_{\odot}$, which is the extreme for which all gas is blown out ($f_{\rm gas}=0$) in halos of this mass. In addition, the changes in $\alpha$ and $M_{\mathrm{crit}}$ lead to a slower decrease in the ejected mass fraction with halo mass, which increases the ejected mass fractions in all halos. However, we ensure that $f_{\rm eject}=0$ for halos with $M_{\rm DMO}\geq 10^{15}M_{\odot}$. We refer to this as the ‘extreme-ejection’ model. We also set the transition radius for the gas profile at $x_{\rm tr}=c_{200}$ (see Appendix A).This configuration induces the strongest baryonic effects on satellite kinematics.
@@
The top panels of Fig. 5 show the cumulative mass fractions of the baryonic components as a function of halo mass for the three models considered here: the ‘no-ejection’ model (left), the fiducial model (middle), and the ‘extreme-ejection’ model (right). In all three cases, the mass fractions of the central galaxy and the diffuse stellar component are identical, the only variation being the ejected mass fraction, $f_{\rm eject}(M_{\rm DMO})$, which leads to different gas fractions. The total bound baryonic fraction is the sum of the mass fractions of the central galaxy, diffuse stars and gas, which is by definition equal to $1-f_{\rm eject}$. In the no-ejection model, the bound baryonic fraction matches the universal baryon fraction across all halo mass scales. In the fiducial model, the bound baryonic fraction is much lower ($\sim$ 0.3) at mass scales of $10^{12}-10^{13}\>{\rm M_{\odot}}$ but rises to match the universal baryon fraction at $M_{\rm DMO}\sim 10^{14}\>{\rm M_{\odot}}$. Finally, the extreme-ejection model has a baryon mass deficit at all scales, reaching the universal baryon fraction only at a halo mass scale of $10^{15}\>{\rm M_{\odot}}$.
@@
Figure 6: Line-of-sight velocity dispersion profiles for halos of $10^{12.5}\>{\rm M_{\odot}}$ (left) and $10^{13.5}\>{\rm M_{\odot}}$ (right), showing the impact of varying the scatter in the stellar mass–halo mass relation and the size–stellar mass relation. The shaded regions indicate the radial ranges used to compute the aperture velocity dispersion, corresponding to two values of $R_{\rm min}$: the fiducial $5\arcsec\arcsec$ fibre collision scale (grey) and $0.1r_{200}$ (yellow). While stellar properties significantly affect the velocity dispersion at small radii, the fiducial aperture, set by the SDSS fibre collision scale, fortunately excludes these central regions, simplifying the baryonic corrections.
@@ 4.5 Bias on dynamical mass inference due to baryonic effects
In Fig. 7, we present the inferred galaxy–halo connection under three scenarios: (1) no baryonic correction (DMO) (blue), similar to Mitra et al. (2024), but without using the LF as additional constraints; (2) using the fiducial baryonic correction model (pink), and (3) using the extreme ejection model (brown). The left panel presents the median relation between central galaxy luminosity and DMO halo mass. Each shaded region indicates the corresponding 68% confidence interval in the inferred $L_{\rm c}-M_{\rm DMO}$ relation. The right panels show the best-fit halo mass distributions at fixed central galaxy luminosity for three cases: $L_{\rm c}=10^{10.8}\,h^{-2}\>{\rm L_{\odot}}$ (top), $L_{\rm c}=10^{10.2}\,h^{-2}\>{\rm L_{\odot}}$ (middle), and $L_{\rm c}=10^{9.6}\,h^{-2}\>{\rm L_{\odot}}$ (bottom).
The differences between the DMO model and the two baryonic models are most pronounced at the low-mass end. This is expected because in the low-mass end $f_{\rm eject}$ is the largest in the baryonic models, leading to the strongest suppression of the enclosed mass profile in individual halos. Thus, for a given $M_{\rm DMO}$, halos in the fiducial feedback scenario exhibit lower velocity dispersions compared to the DMO case. Conversely, for a given observed velocity dispersion, halos in the baryonic scenario require a higher $M_{\rm DMO}$, compared to the DMO case, to match the same satellite kinematics after feedback-driven suppression. As a result, including the baryonic corrections systematically shifts the inferred $M_{\rm DMO}$ to higher values for central galaxies at fixed luminosity. The shift is larger in the maximum feedback scenario, as expected, but the difference relative to the fiducial case is minimal.
Overall, baryonic effects seem to have a fairly negligible impact on the galaxy-halo connection inferred from satellite kinematics for halos above $10^{13}\>{\rm M_{\odot}}$. However, for halos with masses around $10^{12}\>{\rm M_{\odot}}$, comparable to that of the Milky Way, not accounting for baryonic effects causes the inferred median luminosity of centrals to be overestimated by as much as $0.3$ dex. As demonstrated in a forthcoming paper (Mitra et al. in prep), such systematic effects can propagate to significant systematic errors in cosmological inference.
@@ 5 Summary and Conclusion
In a forthcoming paper, Mitra et al. in prep, we demonstrate that such percent level changes in the inferred halo masses can significantly shift the cosmological constraints. This underscores the need to account for baryonic physics when inferring the galaxy–halo connection and deriving cosmological parameters. Moreover, as discussed in Mitra et al. in prep, one can invert this approach by comparing constraints from satellite kinematics corrected for baryonic effects with those from Planck. This comparison opens a potential window for using satellite kinematics data to constrain $f_{\rm eject}$ as a function of halo virial mass, and thus the efficiency of feedback processes associated with galaxy formation.
@@ Acknowledgments
We are grateful to Joop Schaye and Matthieu Schaller for their help with the EAGLE simulations. FvdB is supported by the National Science Foundation (NSF) through grants AST-2307280 and AST-2407063. This work used the following python packages: Matplotlib (Hunter 2007), SciPy (Virtanen et al. 2020), NumPy (van der Walt et al. 2011), halotools (Hearin et al. 2017), AstroPy (Astropy Collaboration et al. 2022) and colossus (Diemer 2018).
@@ A.1 Variations of properties of stars and gas
In the top-left panel, we vary the stellar mass at fixed halo mass by approximately 0.2 dex, reflecting the typical scatter observed in the stellar-halo mass relation, while keeping the fractions $f_{\rm eject}$ and $f_{\rm*,diffuse}$ fixed [^2] With $f_{\rm eject}$ and $f_{\rm*,diffuse}$ held fixed, the maximum increase in stellar mass that still satisfies $f_{\rm\ast,cen}+f_{\rm gas}+f_{\rm eject}+f_{\rm*,diffuse}=1$, and for which $f_{\rm gas}=0$, is only +0.17dex at $M_{\rm DMO}\sim 10^{12}\>{\rm M_{\odot}}$.. As a result, the gas fraction changes accordingly; a higher stellar mass implies more mass has been converted from gas to stars, and vice versa. As is evident, this has no noticeable impact on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. Hence, redistributing mass between gas and stars by an amount that is compatible with the scatter in the SHMR has a negligible effect on the baryonic corrrection. In the top-middle panel, we again vary the stellar mass at a given halo mass by $\pm$0.2 dex, but here we keep the gas fraction ($f_{\rm gas}$) from the fiducial model fixed. As a result, when stars are added or removed, $f_{\rm eject}$ has to change accordingly, meaning we effectively vary $f_{\rm eject}$ here. This has a stronger effect, but one that remains well within the gray zone defined by the more extreme models for $f_{\rm eject}$ presented in Section 4.3. The top right panel shows extreme models for the diffuse stellar component, varying $f_{\rm,diffuse}$ from zero to twice its fiducial value, while keeping $f_{\rm eject}$ fixed. This effectively redistributes gas and diffuse stars in the halo, and has as a negligible effect.
In the middle-left panel, we vary the galaxy size at fixed stellar mass by $\pm$0.5 dex, corresponding to the typical scatter observed in galaxy size-mass relation. This variation again has a negligible influence on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. Likewise, changing the concentration of the diffuse stellar component ($\eta$) by a factor of two produces similarly insignificant effects, as shown in the middle-right panel. In the middle-middle panel, we vary the transition radius of the gas profile to $x_{\rm tr}=0.3\,c_{200}$ and $x_{\rm tr}=c_{200}$. The latter results in a $\sim$0.5% reduction in the velocity dispersion ratio at the high mass end. This outcome is easily understood: for massive halos, where $f_{\rm eject}\rightarrow 0$, the universal baryon fraction—roughly 16% of the total mass—is retained and most of it resides in the gas component. Since the gas profile transitions from a flat core to an NFW-profile, the velocity dispersion of satellite galaxies is lower when this transition occurs further out. We account for this effect in the extreme scenario discussed in Section 4.3, which combines the extreme ejection model, maximal halo response ($\nu=1$), and a larger transition radius ($x_{\rm tr}=c_{200}$).

### v2

@@ 4.2 Comparison with EAGLE
Fig. 3 compares the fiducial model to halos in EAGLE for two different masses: $M_{\rm DMO}=10^{12.5}\>{\rm M_{\odot}}$ (left panel) and $M_{\rm DMO}=10^{13.5}\>{\rm M_{\odot}}$ (right panel). Solid lines show the components of the analytical model, with different colors corresponding to different mass components, as indicated. The shaded regions in matching colors for each component indicate the halo-to-halo variation in EAGLE halos.
@@
Finally, the total density profiles show good agreement between the EAGLE simulations and our fiducial analytical model. This is noteworthy because, although the simulation data was used to calibrate $f_{\rm\ast,diffuse}(M_{\rm DMO})$ and $f_{\rm eject}(M_{\rm DMO})$, the model was not tuned to reproduce any of the density profiles in EAGLE. Since the kinematics of the satellite galaxies are ultimately controlled by the total enclosed mass profile, we can thus be confident that our fiducial analytical model is meaningful.
@@ 4.3 Variations from the fiducial model
Figure 4: Illustrations of the halo response. Left: Solving for the response of the dark matter due to the presence of the baryons for the fiducial model with $\nu=1$ for different halo masses. The final radius $r_{\rm f}$ is plotted against initial radius $r_{\rm i}$. The dashed line shows $r_{\rm f}=r_{\rm i}$, indicating when there would be no halo response. For small $r_{\rm i}$ we see that $r_{\rm f}<r_{\rm i}$, indicating that dark matter that was initially at larger radii migrated inwards (contraction). On the other hand, at at large $r_{\rm i}$, we see that $r_{\rm f}\geq r_{\rm i}$, indicating a very moderate expansion. Both effects become weaker with increasing halo mass. Middle: The line-of-sight velocity dispersion profile for a halo of mass $M_{\rm DMO}\sim 10^{12.5}\>{\rm M_{\odot}}$, for $\nu=0$ (blue), $\nu=0.5$ (red) and $\nu=1$ (green). Under $\nu>0$, we again see that the dark matter halo responds by contracting in the center and expanding in the outskirts, though only the latter affects satellite kinematics data (gray shaded region). Right: Impact of the halo response on the aperture velocity dispersion ratio, $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$, as a function of halo mass. The blue curve indicates no halo response $\nu=0$, equivalent to $r_{\rm i}=r_{\rm f}$. When the dark matter responds ($\nu>0$), the observed effect is a further decrease of $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$.
@@ 4.3.1 Varying halo response strength
Figure 5: Comparison of the effects of the different ejection models. Top panel: The cumulative mass fractions of each baryonic component as a function of halo mass for the three scenarios considered (see text): the ’no-ejection model’ (left), the fiducial model (middle), and the ’extreme-ejection model’ (right). In each scenario, we keep the mass fractions of the central galaxy (red) and diffuse stellar component (pink) fixed, and we only vary the parameter $f_{\rm eject}$ (brown), which leads to different total gas fractions (blue). The total bound baryonic fraction, shown as the black dashed line, is simply obtained by summing all bound baryonic fractions, which is by definition equal to 1 minus the ejected fraction. Bottom panel: the impact of baryonic effects on the ratio of integrated velocity dispersion $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}$ as a function of halo mass. Without baryonic corrections, the ratio remains at 1 (dotted black line). The blue curves, assuming no dark matter halo response, follow roughly $\sigma_{\rm ap}/\sigma_{\rm ap,DMO}\approx$ $\sqrt{M_{\rm tot}/M_{\rm DMO}}=\sqrt{1-f_{\rm b}f_{\rm eject}}$ (dashed black line). The green curves include halo response when preserving adiabatic invariants ($\nu=1$), further reducing velocity dispersion ratio. The gray shaded region represents the range defined by the extreme models, spanning zero to maximal baryonic effect.
@@ 4.3.2 Varying ejection strength
In this section, we examine the impact of changes of the parameters of the function $f_{\rm eject}(M_{\rm DMO})$ given in equation (19). Since mass conservation requires $f_{\rm gas}+f_{\rm*,cen}+f_{\rm*,diffuse}+f_{\rm eject}=1$, such changes also imply adjustment to one or more other mass fractions. To fully understand the consequences of these trade-offs, we test multiple configurations and define two limiting models that represent extreme, yet self-consistent, baryonic configurations.
@@
In the first extreme scenario, we keep $f_{\rm*,cen}$ and $f_{\rm*,diffuse}$ at their fiducial values, but set $f_{\rm eject}=0$. This means that we are not ejecting any baryons (e.g. no stellar or AGN outflows); for halos at all mass scales, the gas now complements the stars such that the total bound baryon fraction adds up to the universal baryon fraction. Since the gas contributes to the mass in the outskirts of the halo, where it has a density profile that is similar to that of the dark matter, this model leads to minimal baryonic effects on satellite kinematics. We refer to this configuration as the ‘no-ejection’ model.
@@
In the second extreme model, we again keep the fiducial fractions for the central galaxy and diffuse stellar component, but we change the gas ejection parameters that describe the mass dependence of $f_{\rm eject}$ to $\alpha=1.2$, $\log M_{\mathrm{char,eject}}=13.7$ and $f_{\mathrm{eject,max}}=0.8$. These values give $f_{\rm eject}\sim 0.8$ at $10^{12}M_{\odot}$, which is the extreme for which all gas is blown out ($f_{\rm gas}=0$) in halos of this mass. In addition, the changes in $\alpha$ and $M_{\mathrm{char,eject}}$ lead to a slower decrease in the ejected mass fraction with halo mass, which increases the ejected mass fractions in all halos. However, we ensure that $f_{\rm eject}=0$ for halos with $M_{\rm DMO}\geq 10^{15}M_{\odot}$. We refer to this as the ‘extreme-ejection’ model. We also set the transition radius for the gas profile at $x_{\rm tr}=c_{200}$ (see Appendix A).This configuration induces the strongest baryonic effects on satellite kinematics.
@@
The top panels of Fig. 5 show the cumulative mass fractions of the baryonic components as a function of halo mass for the three models considered here: the ‘no-ejection’ model (left), the fiducial model (middle), and the ‘extreme-ejection’ model (right). In all three cases, the mass fractions of the central galaxy and the diffuse stellar component are identical, the only variation being the ejected mass fraction, $f_{\rm eject}(M_{\rm DMO})$, which leads to different gas fractions. The total bound baryonic fraction is the sum of the mass fractions of the central galaxy, diffuse stars and gas, which is by definition equal to $1-f_{\rm eject}$. In the no-ejection model, the bound baryonic fraction matches the universal baryon fraction across all halo mass scales. In the fiducial model, in the absence of halo response to baryons ($\nu=0$), the bound baryonic fraction is much lower ($\sim$ 0.3) at mass scales of $10^{12}-10^{13}\>{\rm M_{\odot}}$ but rises to match the universal baryon fraction at $M_{\rm DMO}\sim 10^{14}\>{\rm M_{\odot}}$. Finally, the extreme-ejection model has a baryon mass deficit at all scales, reaching the universal baryon fraction only at a halo mass scale of $10^{15}\>{\rm M_{\odot}}$.
@@
Figure 6: Line-of-sight velocity dispersion profiles for halos of $10^{12.5}\>{\rm M_{\odot}}$ (left) and $10^{13.5}\>{\rm M_{\odot}}$ (right), showing the impact of varying the scatter in the stellar mass–halo mass relation ($\Delta M_{\rm*}$) and the size–stellar mass relation ($\Delta R_{\rm e}$). The solid blue curve shows the fiducial model without halo response ($\nu=0$), while the solid green curve corresponds to the same fiducial baryonic configuration but including adiabatic halo response ($\nu=1$). For all variations in $\Delta M_{\rm*}$ and $\Delta R_{\rm e}$ the halo response parameter $\nu=0$. The shaded regions indicate the radial ranges used to compute the aperture velocity dispersion, corresponding to two values of $R_{\rm min}$: the fiducial $5\arcsec\arcsec$ fibre collision scale (grey) and $0.1r_{200}$ (yellow). While stellar properties significantly affect the velocity dispersion at small radii, the fiducial aperture, set by the SDSS fibre collision scale, fortunately excludes these central regions, simplifying the baryonic corrections.
@@ 4.5 Bias on dynamical mass inference due to baryonic effects
In Fig. 7, we present the inferred galaxy–halo connection under three scenarios: (1) no baryonic correction (DMO) (blue), similar to Mitra et al. (2024), but without using the LF as additional constraints; (2) using the fiducial baryonic correction model (pink), and (3) using the extreme ejection model (brown). The left panel presents the median relation between central galaxy luminosity and DMO halo mass. Each shaded region indicates the corresponding 68% confidence interval in the inferred $L_{\rm c}-M_{\rm DMO}$ relation. The right panels show the best-fit halo mass distributions at fixed central galaxy luminosity for three cases: $L_{\rm c}=10^{10.8}\,h^{-2}\>{\rm L_{\odot}}$ (top), $L_{\rm c}=10^{10.2}\,h^{-2}\>{\rm L_{\odot}}$ (middle), and $L_{\rm c}=10^{9.6}\,h^{-2}\>{\rm L_{\odot}}$ (bottom). We note that for the highest luminosity bin (top right panel), the posterior peaks at a lower mass than indicated by the median relation in the left-hand panel (where $L_{\rm c}\approx 10^{10.8}\,h^{-2}\>{\rm L_{\odot}}$ corresponds to $M_{\rm DMO}\sim 10^{15}\>{\rm M_{\odot}}$). This is not due to a prior on the halo mass function, but simply reflects the exponential suppression of the halo mass function at the high-mass end.
The differences between the DMO model and the two baryonic models are most pronounced at the low-mass end. This is expected, because in the low-mass end $f_{\rm eject}$ is the largest in the baryonic models, leading to the strongest suppression of the enclosed mass profile in individual halos. Thus, for a given $M_{\rm DMO}$, halos in the fiducial feedback scenario exhibit lower velocity dispersions compared to the DMO case. Conversely, for a given observed velocity dispersion, halos in the baryonic scenario require a higher $M_{\rm DMO}$, compared to the DMO case, to match the same satellite kinematics after feedback-driven suppression. As a result, including the baryonic corrections systematically shifts the inferred $M_{\rm DMO}$ to higher values for central galaxies at fixed luminosity. The shift is larger in the maximum feedback scenario, as expected, but the difference relative to the fiducial case is minimal.
Overall, baryonic effects seem to have a fairly negligible impact on the galaxy-halo connection inferred from satellite kinematics for halos above $10^{13}\>{\rm M_{\odot}}$. However, for halos with masses around $10^{12}\>{\rm M_{\odot}}$, comparable to that of the Milky Way, not accounting for baryonic effects causes the inferred median luminosity of centrals to be overestimated by as much as $0.3$ dex. As demonstrated in a subsequent paper (Mitra et al. 2025), such systematic effects can propagate to significant systematic errors in cosmological inference.
@@ 5 Summary and Conclusion
In a companion paper (Mitra et al. 2025), we demonstrate that such percent level changes in the inferred halo masses can significantly shift the cosmological constraints. This underscores the need to account for baryonic physics when inferring the galaxy–halo connection and deriving cosmological parameters. Moreover, as discussed in Mitra et al. (2025), one can invert this approach by comparing constraints from satellite kinematics corrected for baryonic effects with those from Planck. This comparison opens a potential window for using satellite kinematics data to constrain $f_{\rm eject}$ as a function of halo virial mass, and thus the efficiency of feedback processes associated with galaxy formation.
@@ Acknowledgments
We are grateful to Joop Schaye and Matthieu Schaller for their help with the EAGLE simulations. FvdB is supported by the National Science Foundation (NSF) through grants AST-2307280 and AST-2407063. KM’s work at Argonne was supported under the DOE contract DE-AC02-06CH11357. This work used the following python packages: Matplotlib (Hunter 2007), SciPy (Virtanen et al. 2020), NumPy (van der Walt et al. 2011), halotools (Hearin et al. 2017), AstroPy (Astropy Collaboration et al. 2022) and colossus (Diemer 2018). Finally, we are very grateful to the referee, Gary Mamon, for an extremely careful and detailed report that helped strengthen this work.
@@ A.1 Variations of properties of stars and gas
In the top-left panel, we vary the stellar mass of the central galaxy at fixed halo mass by approximately 0.2 dex, reflecting the typical scatter observed in the stellar-halo mass relation, while keeping the fractions $f_{\rm eject}$ and $f_{\rm*,diffuse}$ fixed [^2] With $f_{\rm eject}$ and $f_{\rm*,diffuse}$ held fixed, the maximum increase in stellar mass that still satisfies $f_{\rm\ast,cen}+f_{\rm gas}+f_{\rm eject}+f_{\rm*,diffuse}=1$, and for which $f_{\rm gas}=0$, is only +0.17dex at $M_{\rm DMO}\sim 10^{12}\>{\rm M_{\odot}}$.. As a result, the gas fraction changes accordingly; a higher stellar mass implies more mass has been converted from gas to stars, and vice versa. As is evident, this has no noticeable impact on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. Hence, redistributing mass between gas and stars by an amount that is compatible with the scatter in the SHMR has a negligible effect on the baryonic correction. In the top-middle panel, we again vary the stellar mass of the central galaxy at a given halo mass by $\pm$0.2 dex, but here we keep the gas fraction ($f_{\rm gas}$) from the fiducial model fixed. As a result, when stars are added or removed, $f_{\rm eject}$ has to change accordingly, meaning we effectively vary $f_{\rm eject}$ here. This has a stronger effect, but one that remains well within the gray zone defined by the more extreme models for $f_{\rm eject}$ presented in Section 4.3. The top right panel shows extreme models for the diffuse stellar component, varying $f_{\rm,diffuse}$ from zero to twice its fiducial value, while keeping $f_{\rm eject}$ fixed. This effectively redistributes gas and diffuse stars in the halo and has a negligible effect.
In the middle-left panel, we vary the central galaxy’s size at fixed stellar mass by $\pm$0.5 dex, corresponding to the typical scatter observed in the galaxy size-mass relation. This variation again has a negligible influence on $\sigma_{\rm ap}/\sigma_{\rm app,DMO}$. Likewise, changing the concentration of the diffuse stellar component ($\eta$) by a factor of two produces similarly insignificant effects, as shown in the middle-right panel. In the middle-middle panel, we vary the transition radius of the gas profile to $x_{\rm tr}=0.3\,c_{200}$ and $x_{\rm tr}=c_{200}$. The latter results in a $\sim$0.5% reduction in the velocity dispersion ratio at the high mass end. This outcome is easily understood: for massive halos, where $f_{\rm eject}\rightarrow 0$, the universal baryon fraction—roughly 16% of the total mass—is retained and most of it resides in the gas component. Since the gas profile transitions from a flat core to an NFW-profile, the velocity dispersion of satellite galaxies is lower when this transition occurs further out. We account for this effect in the extreme scenario discussed in Section 4.3, which combines the extreme ejection model, maximal halo response ($\nu=1$), and a larger transition radius ($x_{\rm tr}=c_{200}$).